### v3

# Seele: A Unified Acceleration Framework for Real-Time Gaussian Splatting on Mobile Devices

Abstract
3D Gaussian Splatting (3DGS) has become a crucial rendering technique for many real-time applications. However, the limited hardware resources on today’s mobile platforms hinder these applications, as they struggle to achieve real-time performance. In this paper, we propose Seele, a general framework designed to accelerate the 3DGS pipeline for resource-constrained mobile devices.
Specifically, we propose two GPU-oriented techniques: hybrid preprocessing and contribution-aware rasterization. Hybrid preprocessing alleviates the GPU compute and memory pressure by reducing the number of irrelevant Gaussians during rendering. The key is to combine our view-dependent scene representation with online filtering. Meanwhile, contribution-aware rasterization improves the GPU utilization at the rasterization stage by prioritizing Gaussians with high contributions while reducing computations for those with low contributions. Both techniques can be seamlessly integrated into existing 3DGS pipelines with minimal fine-tuning. Collectively, our framework achieves up to 6.3 speedup and 39.1% model reduction while achieving superior rendering quality compared to existing methods. Our codes will be released upon publication.
1 Introduction
3D Gaussian Splatting (3DGS) [kerbl20233d] has emerged as a vital technique in many rendering-related domains, including autonomous driving [liu2025citygaussian, liu2024citygaussianv2, kerbl2024hierarchical, tancik2022block], augmented and virtual reality (AR/VR) [rojas2023re, chen2023mobilenerf, hu2022efficientnerf, hedman2021baking]. These real-time applications require high rendering performance to ensure seamless quality of services. However, the limited computational resources of mobile platforms often constrain these applications, preventing them from reaching their full potential.
For instance, Nvidia AGX Orin, a leading-edge compute module for modern vehicles, offers only 3.4% of the computational resources of a Nvidia A100 workstation GPU. Similarly, the Snapdragon XR2 chip, the most widely used chip for AR/VR devices, has 4.2% of the computational power compared to Nvidia RTX 4090. On Nvidia’s AGX Orin, 3DGS [kerbl20233d] barely achieves 20 frames-per-second (FPS) on real-world datasets [barron2021mip, knapitsch2017, hedman2018deep], far from the real-time requirement, i.e., 90 FPS of VR [wang2023effect]. An interesting challenge is to achieve real-time on devices with only 3-4% of the resources compared to high-end GPUs.
To achieve this goal, we must understand the major performance bottlenecks in the current 3DGS rendering pipelines. We summarized the bottlenecks into three aspects: computational intensity, rendering inefficiency, and memory budget. This paper proposes a one-stop solution framework, Seele, that addresses the inefficiencies across these three dimensions (Sec. 3). Next, we briefly discuss the inefficiencies in today’s 3DGS rendering pipeline and highlight how our proposed solution addresses these limitations.
Computational Intensity.
One primary challenge in the 3DGS pipeline is its computational intensity. From an individual pixel’s perspective, rendering one pixel typically involves processing thousands of Gaussian points, with each GPU thread rendering one pixel. This rendering process imposes substantial computational overhead on resource-constrained devices. Prior solutions [fan2023lightgaussian, fang2024mini] primarily focus on reducing the overall model size, namely the number of Gaussian points, but these methods often trade off rendering quality for performance.
In contrast, we propose hybrid preprocessing with a view-dependent scene representation that dynamically loads the relevant Gaussians to GPU memory during runtime (Sec. 3.3). Our data representation naturally identifies the necessary Gaussians contributing to the current viewpoint while leaving other irrelevant Gaussians untouched. Meanwhile, our method does not make any compromise between performance and accuracy. On average, this method alone outperforms the state-of-the-art rendering algorithms while achieving a 2.8 speedup.
Rendering Inefficiency.
The second issue with the current 3DGS rendering pipeline is that all splatted Gaussians go through the same rasterization pipeline regardless of their contribution to the final pixels. Our experiment shows that, for each pixel, 1.5% of the Gaussians contribute to 99% of the final pixel. Therefore, we argue that this uniform treatment to all Gaussians leads to severe rendering inefficiencies, as less significant Gaussians should naturally receive fewer computational budgets. However, prior work [lee2024compact, girish2025eagles, lin2024rtgs], such as pruning or quantization techniques, still retains a uniform rendering pipeline and fails to differentiate Gaussians based on their contributions.
On the contrary, we propose contribution-aware rasterization from a pixel-centric perspective (Sec. 3.4). Rather than reducing computation uniformly, we dynamically identify and prioritize Gaussians with high contributions (i.e., high accumulative transparencies) and reduce computation budgets for those with low contributions. Gaussians with low contributions, which typically contribute to low-frequency textures (as shown in Fig. 4), are allocated less computation by dynamically skipping insignificant operations. We show that this method achieves an additional 1.4 speedup while retaining the same rendering quality.
Memory Budget.
Our last contribution eases the peak GPU memory. As model complexity scales up, the GPU memory requirements often become unsustainable due to the increasing number of Gaussians involved. Solutions [lee2024compact, ren2024octree] apply offline compression techniques to reduce memory usage, but they overlook a key opportunity: rendering a given viewpoint does not require loading all Gaussians into GPU memory simultaneously.
Leveraging this key observation, we design proactive memory management that asynchronously loads necessary data to GPU memory without disrupting the critical path of rendering. Combined with our view-dependent scene representation, our memory management significantly reduces peak GPU memory usage. On average, our approach achieves a 39.1% reduction in model size, allowing large-scale rendering on mobile devices.
Collectively, our framework achieves up to 6.3 speedup and 39.1% runtime model reductions, all while retaining a better rendering quality compared to state-of-the-art algorithms. Furthermore, our techniques are orthogonal to existing optimizations and can be seamlessly integrated with other mainstream Gaussian splatting pipelines.
The contribution of this paper is summarized as follows:
We introduce a view-dependent scene representation that combines online and offline filtering to reduce computation overhead and runtime GPU memory.
We propose a contribution-aware rasterization algorithm that dynamically skips the insignificant computations and improves the parallel efficiency.
We design an integrated co-training procedure that integrates the aforementioned techniques and achieves better rendering quality against the corresponding baselines.
2 Related Work
2.1 Efficient Data Representation
A 3DGS model is essentially a set of points in 3D space that represent the physical structure of the world. The model size often becomes one of the key bottlenecks in rendering performance and memory usage. To reduce the model size, several studies have proposed various data representations to optimize storage and performance. Some [niedermayr2024compressed, chen2024hac, chen2025hac++, papantonakis2024reducing, durvasula2025contrags, niemeyer2024radsplat] reduce GPU memory usage by leveraging encoding techniques to compress the model. For example, CompactGS [lee2024compact] employs vector quantization to compress the model offline and decodes it at runtime using a pretrained codebook. Similarly, EAGLES [girish2025eagles] uses an encoding network to compress the neural radiance field into a compact representation. However, both methods introduce non-trivial execution overhead during runtime, which limits their applicability in computation-constrained scenarios.
Other pruning-based approaches use a learnable mask [liu2025maskgaussian, zhang2024lp, taghipour2025svr] or a significance score [fang2024mini, fang2024mini2, mallick2024taming, fan2023lightgaussian, hanson2025pup] to eliminate redundant Gaussians. However, these methods have to make a trade-off between rendering quality and efficiency. Some studies adopt structured data representations such as octrees [ren2024octree] and kd-tree [kerbl2024hierarchical] to organize Gaussian points while some researches [hamdi2024ges, held20253d, li20253d, ye2025gaussian] explore alternative primitives for more efficient scene representation. However, they introduce additional runtime overhead in loading or rasterizing Gaussians.
In contrast to prior studies, we focus on real-time performance. Our view-dependent scene representation achieves significant model size reduction with minimal runtime overhead and no quality compromise.
2.2 Rasterization Optimization
In addition to pruning Gaussian points, another line of research focuses on optimizing the performance of the 3DGS pipeline itself. Recent studies propose various online filtering techniques, such as axis-aligned bounding box (AABB) [klosowski1998efficient, wang2024adr, hanson2025speedy] and oriented bounding box (OBB) [gottschalk1996obbtree, lee2024gscore] intersection tests, to perform more fine-grained filtering before rasterization. These works aim to reduce the number of Gaussians processed during the rasterization stage at runtime, thereby reducing the computational workload.
On the other hand, some methods [feng2025flashgs, liao2025tc, gui2024balanced] focus on optimizing the rasterization stage. For example, FlashGS [feng2025flashgs] designs a software pipelining to overlap data fetching with rasterization, improving overall efficiency. In contrast, Balanced3DGS [gui2024balanced] addresses workload imbalance by introducing an offline scheduling mechanism to rebalance workloads at both the block and tile levels. Additionally, a few studies [lee2024gscore, lin2025metasapiens, pei2025gcc, lee2025vr] explore the new hardware designs to support 3DGS, further advancing the efficiency and scalability of the pipeline.
Unlike prior works, our optimization takes a unique angle to accelerate 3DGS by addressing the inherent contribution imbalance across Gaussian points. Our method dynamically assigns computational budgets to Gaussians based on their contributions to the final rendered image. Moreover, our algorithm is naturally GPU-friendly and alleviates execution inefficiencies such as warp divergence, achieving better overall performance with no hardware modifications.
3 Methodology
3.1 Preliminaries
3DGS algorithms adhere to an explicit point-based representation, where each point is modeled by a Gaussian ellipsoid (a.k.a., Gaussian for short) that captures the geometric and textural properties of the scene. The geometric property of a Gaussian is governed by its 3D centroid position and a 3D covariance matrix . The textural properties are described by two components: its opacity and spherical harmonic (SH) coefficients. Using these properties, the 3DGS pipeline renders a frame on a tile-by-tile basis via three steps: preprocessing, sorting, and rasterization.
Preprocessing.
Given a camera pose and a view direction, the preprocessing stage first filters out Gaussians outside the current view frustum. Meanwhile, this stage identifies the intersection between Gaussians and rendering tiles.
Sorting.
Once each tile collects its intersected Gaussians, the sorting stage determines the rendering order of those Gaussians. This stage ensures that all points are rendered, from the closest to the furthest, based on their depth.
Rasterization.
Once all Gaussians are sorted, the rasterization stage renders these Gaussians tile-by-tile. Within each tile, every pixel iterates through the same set of Gaussians, computing transparency and accumulating their colors into the pixel in the sorted order. Eqn. 1 governs the color accumulation process of pixel p,
where denotes the accumulative transmittance of pixel p from the Gaussian st to the th. and stand for the transparency and the color at the th Gaussian, respectively. The transparency of a Gaussian is calculated by the Gaussian opacity , the 2D position and the 2D covariance matrix, [zwicker2001ewa],
is a function of the Gaussian spherical harmonics.
Note that, if the Gaussian’s value is insignificant (less than , i.e., ), this Gaussian will be skipped in the color accumulation to avoid numerical instabilities. The color accumulation process terminates once the cumulative transmittance surpasses a predefined threshold, .
3.2 Overview
Fig. 2 gives the overview of our proposed acceleration framework. To address three rendering inefficiencies highlighted in Sec. 1, we introduce two key components: hybrid preprocessing and contribution-aware rasterization to reduce the runtime computation and peak memory.
In 3DGS pipelines, rasterization is the most time-consuming stage, accounting for 67% of the execution time on a Nvidia Ampere GPU. Prior studies have explored two primary methods to reduce the overall workload of rasterization. The first approach applies pruning techniques, which eliminate insignificant Gaussians offline [fan2023lightgaussian, fang2024mini]. The second proposes online filtering techniques, such as AABB and OBB intersection tests, to perform more fine-grained filtering before rasterization [klosowski1998efficient, wang2024adr, gottschalk1996obbtree, lee2024gscore].
Our hybrid preprocessing combines the merits of both offline and online techniques. During the offline processing, we design a view-dependent scene representation that clusters Gaussians based on their contributions to a set of closely related rendering viewpoints. At runtime, our framework can asynchronously prefetch the relevant clusters into the GPU memory before rendering. Our view-dependent scene representation inherently eliminates Gaussians that are irrelevant to the current rendering while within the view frustum. Combined with online filtering, we further reduce the number of Gaussians intersected for each tile in the subsequent stages.
Meanwhile, we propose a novel contribution-aware rasterization to accelerate the rasterization pipeline itself. The key idea of our approach is to penalize the computations assigned to insignificant Gaussians purposely, allowing for efficient rendering without compromising quality.
Specifically, our rasterization pipeline organizes pixels into small groups while still keeping one thread responsible for one pixel. At runtime, only a single pixel within each pixel group calculates its Gaussian transparency rather than having all pixels compute . The purpose is to let this pixel identify Gaussians that potentially contribute less to the pixel group, so we can exclude these Gaussians from further rasterization. In this way, we eliminate insignificant Gaussian accumulation at runtime. Combined with our integrated co-training, we can achieve a smoother rendering experience. A more rigorous explanation of this approach is provided in Sec. 3.4.
3.3 Hybrid Preprocessing, HP
View-Dependent Scene Representation.
The idea of this data representation is to preprocess Gaussians offline, thereby reducing computational effort during the online preprocessing stage. Our key observation is that different viewpoints rely on distinct sets of Gaussians for rendering. Viewpoints in close proximity are likely to share similar Gaussians, while those farther apart share fewer.
Based on this observation, our data representation classifies Gaussians into two categories: shared Gaussians and exclusive Gaussians, as shown in Fig. 3. Shared Gaussians are shared across all viewpoints and are frequently used for rendering frames from any viewpoint (yellow points in Fig. 3(b)). On the other hand, exclusive Gaussians are specifically attached to a group of viewpoints and are only used for rendering frames within that group (other colors in Fig. 3(b)).
Offline Clustering.
The key to our approach lies in how to classify the Gaussians into shared and exclusive Gaussians. Initially, our method randomly samples camera poses and clusters them into smaller groups based on the weighted similarity of their camera positions, and view directions, . Specifically, we first normalize the camera positions, , then concatenate normalized and as a 1D dimensional vector, , where is a hyperparameter that balances the impact of position and orientation. Here, is set to 1 for clustering all Gaussians.
For each cluster, we identify and reserve the Gaussians that are the primary contributors to pixel rendering. Specifically, for each pixel, we find the Top- Gaussians based on their contribution to the cumulative transmittance, , as the contribution metric. The rationale here is that the remaining Gaussians contribute minimally to the final pixel and can be safely discarded without compromising rendering quality. We choose a value of 32, as we find that increasing further yields minimal quality improvement.
Once we identify the top Gaussian contributors for each cluster, we find the union of top contributors across all clusters and grant these top Gaussians to be shared across all clusters. These granted Gaussians are then the shared Gaussians in our representation. The remaining Gaussians within each cluster are classified as exclusive Gaussians.
Algorithm.
During the actual rendering, for a given camera pose, our approach first identifies the nearest cluster based on the cluster centroid. Next, all Gaussians belonging to this cluster and its neighboring clusters are selected, as shown in Fig. 2. In this example, green points are the nearest cluster, and the yellow points are the nearest neighbors to that cluster ( is set to 1 in this toy example). The selected clusters are used to render the current camera pose instead of all Gaussians. Note that, shared Gaussians, gray points in Fig. 2, are always stored in GPU memory since they are used for any poses. Unlike the original 3DGS, our approach retains only the relevant clusters to stay in GPU memory, thus reducing the peak GPU memory usage. Empirically, we set to be 3, which strikes a good balance between performance and rendering quality.
Online Filtering.
Once these Gaussians are loaded, they are transformed into screen space with a new 2D Gaussian position and covariance matrix . To filter unnecessary Gaussians for each tile, the intersection test typically uses envelope to approximate,
where is the pixel position. Due to the redundancy of this approximation [hanson2025speedy, pei2025gcc], we further include the opacity into our online filtering equation as follows. As we would skip insignificant , which value is less than , , Eqn. 2 can be expressed as,
Therefore, we execute fine-grained online filtering by including opacity into consideration,
This further eliminates false-positive intersected Gaussians when performing Gaussian-tile intersections.
Scene Prefetching.
Our hybrid preprocessing not only enhances rendering performance but also reduces peak GPU memory usage. However, loading Gaussian clusters at runtime could introduce non-trivial execution overhead and potentially cause frame stuttering. To address this, we prefetch future clusters ahead of time by asynchronously loading them using a dedicated GPU stream. We adopt similar prediction methods as prior works [hou2020motion, feng2024cicero] and find that linearly extrapolating the camera pose is an effective method for predicting future poses in our scene prefetching. This way, our approach enables asynchronous prefetching to be overlapped with runtime frame rendering. Compared to loading all Gaussian clusters into GPU memory, our scene prefetching reduces the runtime model size by 39.1% with a minimal runtime overhead (6% of the total latency).
3.4 Contribution-Aware Rasterization, CR
Motivation.
The goal of our contribution-aware rasterization is to pay “computational” attention to significant Gaussians that contribute more to the final pixel values while reducing computations for less important Gaussians. Our observation is that the color values of most pixels are determined by a small fraction of Gaussians, as illustrated in Fig. 4. We characterize the MipNeRF360 dataset [barron2021mip] and draw the cumulative transparency of each pixel by sorting the Gaussian transparency contribution in descending order. The result in Fig. 4 shows that 1.5% of the top Gaussians contribute to 99% of the final pixel, while remaining Gaussians contribute minimally due to the lower sampling . Furthermore, these low contribution samples typically occur in Gaussians with larger variances or at locations farther from the mean, corresponding to the low-frequency components and sharing low gradients and good spatial locality. More rigorous analysis of the correlation between Gaussians and frequency is shown in the supplementary.
Algorithm.
To leverage the observation, we propose a contribution-aware rasterization algorithm to identify insignificant samples and rearrange computational resources, as outlined in Algo. 1. Instead of uniformly assigning the same computation to all Gaussians, our algorithm organizes every pixels into a small group, P. During each iteration, only the leader pixel within a group, which is the centroid of this group, computes the value for a given Gaussian, . If this value falls below a predefined threshold (, to avoid numerical instability), the algorithm skips the color blending of this Gaussian on all pixels within P. We set to 2 to balance rendering efficiency and quality, and analyze the upper bound on the blending error introduced by this skipping in the appendix.
For Gaussians deemed important (i.e., those with values above the threshold, ), our algorithm performs the same rasterization computations as the original approach, ensuring rendering quality is preserved. This dynamic computation reassignment improves rendering efficiency without compromising visual quality.
GPU Parallelism.
Not only does our algorithm reduce the overall computation, but it also inherently achieves higher parallelism by alleviating warp divergence. In the GPU implementation, pixels within a group are assigned to one warp. All pixels within a warp execute in “lockstep” as shown in Fig. 5. For instance, when blending Gaussian , only thread accumulates the insignificant Gaussian into its pixel, while other threads wait for thread to complete.
In contrast, our algorithm only lets thread detect if Gaussian is insignificant. In this case, all threads in this warp skip the color blending of Gaussian . In the case of Gaussian , if it is considered significant, all threads would blend Gaussian . This ensures that these pixels either skip or process the same Gaussians collectively. This way, we eliminate divergence within a warp.
3.5 Integrated Fine-Tuning
Note that, both HP and CR modify the rendering pipeline and might impact the view consistency. To maintain a better view consistency across frames, our algorithm only requires minimal fine-tuning to achieve better rendering quality. Since both hybrid preprocessing and contribution-aware rasterization do not participate in gradient descent, these two techniques can be integrated into the current training process without any modifications to back-propagation.
Once the 3DGS model is trained by the canonical training process, we convert the model into our view-dependent scene representation in Sec. 3.3. Next, we fine-tune the model for additional 1000 steps with our loss function,
where is the original loss function in 3DGS and is the view consistency loss. Here, we use F LIP score [andersson2020flip] with 7 frames as our metric. is set to be 0.1.
and
denote the best and second-best results among all methods, respectively.| Dataset | Mip-NeRF360 | Tanks&Temples | Deep Blending | ||||||||||||||||||||||||||
| Metrics | Quality | Efficiency | Quality | Efficiency | Quality | Efficiency | |||||||||||||||||||||||
| PSNR | SSIM | LPIPS | FPS |
PSNR | SSIM | LPIPS | FPS |
PSNR | SSIM | LPIPS | FPS |
| 3DGS [kerbl20233d] | 27.46 | 0.812 | 0.223 | 20.79 | 2168.68 | 710.6 | 23.75 |
0.848 |
0.176 | 41.97 | 1034.37 | 430.9 | 29.59 | 0.902 | 0.244 | 23.94 | 1645.11 | 639.9 | |||||||||||
| Seele + 3DGS |
27.72
0.814
0.217
59.67 | 778.25 | 380.9 |
24.02
0.852
0.168
127.80 | 356.15 | 207.6 | 29.79 | 0.903 |
0.240
86.58 | 421.74 | 400.6 | |||||||||||
| MiniSplatting [fang2024mini] | 27.23 | 0.814 | 0.222 | 71.31 | 797.96 | 145.7 | 23.18 | 0.837 | 0.187 | 143.27 | 329.69 | 83.3 |
30.04
0.908
0.243 | 120.01 | 382.86 | 154.9 | |||||||||||
| Seele + MiniSplatting |
27.70
0.822
0.212
131.62
436.41
106.5 | 23.74 | 0.846 |
0.179
268.03
172.40
60.1 |
30.02 |
0.908
0.242
200.62
224.58
129.6 | |||||||||||
| LightGaussian [fan2023lightgaussian] | 27.44 | 0.807 | 0.235 | 30.89 | 1533.50 |
59.4 |
23.82 | 0.842 | 0.189 | 65.60 | 699.71 |
33.8 |
29.74 | 0.901 | 0.250 | 43.17 | 1097.82 |
52.4 |
| Seele + LightGaussian | 27.56 | 0.810 | 0.229 | 76.36 | 589.85 |
39.1
23.91
0.846 | 0.181 | 183.86 | 259.78 |
20.7
29.73 | 0.900 | 0.249 | 131.16 | 324.25 |
41.4
| AdR-Gaussian [wang2024adr] | 26.92 | 0.786 | 0.267 | 65.95 | 885.27 | 288.6 | 23.42 | 0.835 | 0.202 | 114.18 | 442.99 | 203.0 | 29.68 | 0.901 | 0.255 | 78.87 | 485.00 | 330.5 | |||||||||||
| Seele + AdR-Gaussian | 27.19 | 0.790 | 0.261 |
96.25
327.26
147.6 | 23.78 | 0.838 | 0.196 |
193.83
154.87
106.2 | 29.64 | 0.899 | 0.254 |
141.87
178.16
127.7 | |||||||||||
4 Evaluation
4.1 Experimental Setup
Code Optimization.
In addition to the two proposed techniques in Sec. 3, we also implement the following optimizations. First, we apply fine-grained data fetching when loading Gaussian attributes to better utilize shared memory on GPU. Second, we leverage thread-level parallelism to overlap the execution of fetching Gaussian attributes and computation of alpha blending. Lastly, we replace the compute-intensive exp() function with CUDA intrinsic function __exp(), which generates fewer machine instructions once compiled. Note that, all optimizations above are nothing more than engineering hacks to improve the program’s efficiency; we do not claim them as our major contributions.
Datasets.
To show the efficiency and robustness of Seele, we evaluate on three datasets: Mip-NeRF360 [barron2021mip], Tank&Temple [knapitsch2017], and DeepBlending [hedman2018deep]. For quality evaluation, we use PSNR, SSIM, and LPIPS. We also report FPS, the number of executed instructions (#Inst.), and the runtime model size as our performance metrics.
Baselines.
To show the general applicability of our optimization technique, we apply our framework to four state-of-the-art 3DGS algorithms: 3DGS [kerbl20233d], MiniSplatting [fang2024mini], LightGaussian [fan2023lightgaussian], and AdR-Gaussian [wang2024adr].
Algorithm Configurations.
Hardware.
The primary hardware we used is a mobile Ampere GPU on Nvidia AGX Orin SoC, which is the flagship development board in AR/VR and autonomous driving. Meanwhile, we also report the performance numbers on Nvidia Orin NX, a low-power embedded computing board, and Nvidia A6000, a powerful workstation GPU.
4.2 Performance and Quality
| 3DGS | MiniSplatting | LightGaussian | AdR-Gaussian | |||||
F |
F |
F |
F |
F |
F |
F |
F |
| Baseline | 0.0164 | 0.0466 | 0.0160 | 0.0460 | 0.0168 | 0.0472 | 0.0106 | 0.0332 |
| Seele | 0.0126 | 0.0292 | 0.0110 | 0.0300 | 0.0111 | 0.0295 | 0.0061 | 0.0169 |
Quality Evaluation.
Tbl. 1 shows the overall evaluation of Seele on four widely-used 3DGS pipelines: 3DGS [kerbl20233d], MiniSplatting [fang2024mini], LightGaussian [fan2023lightgaussian], and AdR-Gaussian [wang2024adr]. Generally, Seele achieves better rendering quality on all three datasets across all three evaluation metrics. For instance, on average, Seele improves PSNR and SSIM by 0.28 dB and 0.004, respectively. More qualitative comparisons are shown in the supplementary.
We also measure the view consistency metrics, F LIP1 and F LIP7 [andersson2020flip] in Tbl. 2. With the loss function proposed in Sec. 3.5, Seele achieves better view consistency against the corresponding baselines.
| Dataset | Mip-NeRF360 | Tanks&Temples | Deep Blending | |||
| GPU | Orin-NX | A6000 | Orin-NX | A6000 | Orin-NX | A6000 |
| 3DGS | 3.06 | 134.45 | 6.03 | 197.55 | 3.64 | 155.87 |
| Seele + 3DGS | 9.42 | 328.98 | 19.63 | 536.77 | 14.82 | 515.60 |
| MiniSplatting | 9.62 | 414.02 | 20.57 | 440.34 | 16.84 | 701.51 |
| Seele + MiniSplatting | 17.48 | 640.16 | 41.78 | 1108.65 | 32.40 | 1031.46 |
| LightGaussian | 4.39 | 185.95 | 9.02 | 323.47 | 5.78 | 234.36 |
| Seele + LightGaussian | 12.72 | 431.99 | 27.45 | 755.29 | 20.38 | 656.60 |
| AdR-Gaussian | 10.30 | 383.35 | 16.65 | 456.12 | 13.43 | 478.45 |
| Seele + AdR-Gaussian | 16.70 | 524.88 | 31.84 | 762.71 | 19.57 | 762.45 |
Performance Evaluation.
With comparable rendering quality, Seele consistently achieves speedups across all algorithms. On average, Seele achieves 3.2, 1.8, 2.7, and 1.7 speedup on 3DGS, MiniSplatting, LightGaussian, and AdR-Gaussian respectively. The higher speedup observed in 3DGS compared to the other algorithms can be attributed to the denser nature of 3DGS. This density allows Seele to better separate irrelevant Gaussians contributed to different views using hybrid preprocessing introduced in Sec. 3.3, resulting in greater speedup. Despite that, sparse models like MiniSplatting still benefit from hybrid preprocessing due to the view-dependent redundancy, i.e., different views require different Gaussians to render. Tbl. 1 also shows the average number of executed instructions of each algorithm using NVIDIA Nsight Compute, Seele indeed reduces the executed instructions, thus overall computation. Tbl. 4 further dissects the contributions of our techniques.
Tbl. 3 shows the speedup of Seele on additional GPUs, Nvidia Orin NX and Nvidia A6000. The results show that the optimizations proposed in Seele are not tied to a specific GPU architecture, exhibiting good generality across different hardware platforms. However, we observe that our optimization achieves a higher speedup on lower-power GPUs, as shown in Tbl. 3. Primarily, low-end GPUs often have more restricted hardware resources.
Seele not only speeds up the overall rendering process but also reduces overall GPU memory consumption. Here, we exclusively focus on the GPU memory contributed by the Gaussian points, i.e., runtime model weights. Overall, results show that Seele achieves 39.1% reduction in runtime model size across three different datasets. Even for sparse models like MiniSplatting, Seele still saves 23.2% of model size compared to the baselines.
4.3 Ablation Study
| Dataset | Mip-NeRF360 | Tanks&Temples | Deep Blending | |||||||||||||||
| Metrics | Quality | Efficiency | Quality | Efficiency | Quality | Efficiency | ||||||||||||
| PSNR | SSIM | LPIPS | FPS | #Inst.() | Mem.(MB) | PSNR | SSIM | LPIPS | FPS | #Inst.() | Mem.(MB) | PSNR | SSIM | LPIPS | FPS | #Inst.() | Mem.(MB) | |
| 3DGS [kerbl20233d] | 27.46 | 0.812 | 0.223 | 20.79 | 2168.68 | 710.6 | 23.75 | 0.848 | 0.176 | 41.97 | 1034.37 | 430.9 | 29.59 | 0.902 | 0.244 | 23.94 | 1645.11 | 639.9 |
| 3DGS+Opti. | 27.46 | 0.812 | 0.223 | 21.75 | 1945.88 | 710.6 | 23.75 | 0.847 | 0.176 | 45.96 | 923.51 | 430.9 | 29.59 | 0.902 | 0.244 | 29.60 | 1485.44 | 639.9 |
| 3DGS+Opti.+HP | 27.70 | 0.814 | 0.219 | 46.15 | 814.74 | 380.9 | 24.05 | 0.852 | 0.171 | 118.61 | 374.63 | 207.6 | 29.74 | 0.903 | 0.242 | 81.73 | 412.50 | 400.6 |
| 3DGS+Opti.+CR | 27.50 | 0.812 | 0.221 | 30.10 | 1574.92 | 710.6 | 23.70 | 0.848 | 0.173 | 57.03 | 763.83 | 430.9 | 29.70 | 0.902 | 0.242 | 38.79 | 1156.46 | 639.9 |
| Seele + 3DGS | 27.72 | 0.814 | 0.217 | 59.67 | 778.25 | 380.9 | 24.02 | 0.852 | 0.168 | 127.80 | 356.15 | 207.6 | 29.79 | 0.903 | 0.240 | 86.58 | 421.74 | 400.6 |
In this section, we show the contributions of individual optimizations. Specifically, we evaluate four variants:
+Opti.: this variant only includes the additional code optimizations proposed in Sec. 4.1.
Seele: this variant is the full-fledged algorithm.
Tbl. 4 presents the results of our ablation study on 3DGS. Across all datasets, our code optimization (+Opti.) achieves a 1.1 speedup. Building on this, +Opti.+HP and +Opti.+CR yield additional 2.8 and 1.3 speedups, respectively. When all optimizations are combined, Seele achieves an overall 3.2 speedup. All model weight savings come from HP, as the other two optimizations do not contribute to the model reduction. Note that, Seele may increase the overall instruction count but improves the overall speedup due to the reductions on warp divergence.
In terms of the rendering quality, the additional optimizations proposed in Sec. 4.1 do not change the accuracy. The HP technique from Sec. 3.3, combined with fine-tuning as described in Sec. 3.5, further enhances rendering quality across all three quality metrics. On average, HP improves the rendering quality by 0.23 on PSNR. The effect of CR from Sec. 3.4 on quality is minimal, 0.03 in PSNR.
| Dataset | Mip-NeRF360 | |||||
| Metric | PSNR | SSIM | LPIPS | FPS | #Inst.() | Mem. (MB) |
| w/o Fine-tuning | 27.48 | 0.812 | 0.221 | 59.72 | 768.76 | 380.9 |
| w/ Fine-tuning | 27.72 | 0.814 | 0.217 | 59.67 | 778.25 | 380.9 |
| Dataset | Mip-NeRF360 | Tanks&Temples | Deep Blending | |||
| w/ Loss | w/o Loss | w/ Loss | w/o Loss | w/ Loss | w/o Loss | |
| 3DGS | 0.058 | 0.060 | 0.067 | 0.074 | 0.042 | 0.043 |
Fine-Tuning.
Next, we present the results with and without the integrated fine-tuning, as described in Sec. 3.5. Here, we focus on the results of the Mip-NeRF360 dataset with 3DGS. Similar trends are observed across other datasets in the supplementary. Tbl. 5 shows that fine-tuning results in significant improvements. Meanwhile, fine-tuning does not affect performance for both FPS and GPU memory consumption. Tbl. 6 also shows the effectiveness of , finetuning with our proposed loss term can improve the rendering quality.
4.4 Sensitivity Study
Number of Clusters.
Fig. 6 shows how rendering quality and performance vary with two key hyperparameters, the number of clusters and the number of neighboring clusters during rendering (), in Sec. 3.3. Performance numbers are normalized to our default setting of 24 clusters with 4 neighbors (including itself). As the number of neighboring clusters increases, speedup decreases, while accuracy initially improves. However, too many clusters, e.g., 8 clusters with 8 neighbors, degrades accuracy. This degradation is because each Gaussian is responsible for too many view directions, negatively impacting rendering quality.
Pixel Group.
Fig. 8 illustrates the sensitivity of both the rendering quality and performance to the pixel group size using Mip-NeRF360 [barron2021mip], as defined in Sec. 3.4. Other datasets show a similar trend. As the pixel group size increases, rendering quality initially decreases slightly, but the drop becomes more significant once the group size exceeds . By trading off between quality and performance, we ultimately select the pixel group size of .
Number of Contributors.
Fig. 8 shows the sensitivity of the rendering quality and performance to the number of primary contributors in the clustering strategy (Sec. 3.3). The rendering quality is not significantly affected by the number of contributors, while the performance gradually decreases as the number of contributors increases.
5 Conclusion
Recent neural rendering has transformed the landscape of real-time rendering. 3DGS emerges as a new rendering primitive that offers unprecedented realism using the conventional rasterization pipeline. This paper proposes two principled and GPU-oriented optimizations that drastically improve the general Gaussian splatting pipelines by understanding and leveraging the GPUs’ characteristics. We believe that this work could be the first step to understanding the fundamental limitations of GPUs in supporting 3DGS. The result would be valuable to rethink the next architectures tailored to this future technique.